# Enhancing Quantum Memory Lifetime with Measurement-Free Local Error Correction and Reinforcement Learning

### v1

@@ Appendix G LEC circuits with large gate error rate
On the other hand, as shown in Fig. 16, we observe that the RL-optimized LEC circuit trained at $p_{gate,train}$ performs as much as — or even worse than — the Toom LEC circuit when benchmark $p_{gate}$ is $p_{gate,train}$ if $p_{gate,train}$ is large enough. In other words, we observe that $p_{gate,c}\approx p_{gate,train}$ or even $p_{gate,c}<p_{gate,train}$ for large $p_{gate,train}$ case, unlike as shown in Fig. 11. This feature is the limitation of our RL framework, and this is the reason why we do not investigate improving the threshold gate error rate of Toom’s rule decoder using our RL framework.

### v2

@@ E.4 LEC circuits with large gate error rate
On the other hand, as shown in Fig. 16, we observe that the RL-optimized LEC circuit trained at $p_{\rm{gate,train}}$ performs as much as — or even worse than — the Toom LEC circuit when benchmark $p_{\rm{gate}}$ is $p_{\rm{gate,train}}$ if $p_{\rm{gate,train}}$ is large enough. In other words, we observe that $p_{\rm{gate,c}}\approx p_{\rm{gate,train}}$ or even $p_{\rm{gate},c}<p_{\rm{gate,train}}$ for large $p_{\rm{gate,train}}$ case, unlike as shown in Fig. 10. This feature is the limitation of our RL framework, and this is the reason why we do not investigate improving the threshold gate error rate of Toom’s rule decoder using our RL framework.
Appendix F Details on LEC Application
Figure 17: Method of finding minimum number of global decoding for original vs. hybrid decoding protocol to achieve $P_{\rm{L}}\ <15$% for fixed $T=2000$ with 5000 samples of 2D toric code lattice with $L=16$.
From our choice of the error model, fidelity of CNOT gate $F_{2\rm{Q}}=1-4\times p_{\rm{gate}}$, and fidelity of CCX/CCZ gate $F_{3\rm{Q}}=1-6\times p_{\rm{gate}}$. Thus, multi-qubit gates with $p_{\rm{gate}}=1\times 10^{-4}$ have $F_{\rm{2Q}}=99.96\%$ and $F_{\rm{3Q}}=99.94\%$. In this regime, we investigate what is the minimum number of global MWPM decoding for original vs. hybrid decoding protocol to achieve the final logical error rate after total time $T$ to be below $15\%$.
As shown in Fig. 17, we perform 1D and 2D optimization for the original and hybrid decoding protocol, respectively. For the original decoding, we vary $N_{\rm{MWPM}}$ and check the final $P_{\rm{L}}$ after time $T$. On the other hand, for the hybrid decoding, we first fix a total number of global decoding and then find $N_{\rm{LEC}}$ that minimizes $P_{\rm{L}}$ after time $T$. Then, we choose the minimum number of global MWPM decoding for each of the protocols that satisfy $P_{\rm{L}}<15\%$.
On top of the minimum $N_{\rm{MWPM}}$ for $p_{\rm{gate}}=1\times 10^{-4}$ discussed in Section VI.1, we also compare minimum $N_{\rm{MWPM}}$ for larger $p_{\rm{gate}}$ as well. When $p_{\rm{gate}}=3\times 10^{-3}$, i.e. $F_{2\rm{Q}}=99.88\%$ and $F_{3\rm{Q}}=99.82\%$, we obtain that the hybrid protocol can reduce the number of global decoding by 1 or 2 compared to the original protocol regardless of $T$. Also, note that we observe the same results for these comparisons, even though the original decoding protocol is improved by replacing the 3D MWPM following repeated faulty syndrome extractions with the 2D MWPM following a perfect syndrome extraction.